The length of a rectangle is 6 cm less than twice its width. Find the dimensions of the rectangle if its area is 108 cm.

Answer by checkley77(12844)

LW=108
L=2W-6
(2W-6)W=108
2W^2-6W=108
2W^2-6W-108=0
2(W^2-3W-54)=0
2(W-9)(W+6)=0
W-9=0
W=9 ANS
L=2*9-6
L=18-6
L=12 ANS.
PROOF:
9*12=108
108=108